### v1

# Exploring Cultural Variations in Moral Judgments with Large Language Models

Abstract
Large Language Models (LLMs) have shown strong performance across many tasks, but their ability to capture culturally diverse moral values remains unclear. In this paper, we examine whether LLMs can mirror variations in moral attitudes reported by two major cross-cultural surveys: the World Values Survey and the PEW Research Center’s Global Attitudes Survey. We compare smaller, monolingual, and multilingual models (GPT-2, OPT, BLOOMZ, and Qwen) with more recent instruction-tuned models (GPT-4o, GPT-4o-mini, Gemma-2-9b-it, and Llama-3.3-70B-Instruct). Using log-probability-based moral justifiability scores, we correlate each model’s outputs with survey data covering a broad set of ethical topics. Our results show that many earlier or smaller models often produce near-zero or negative correlations with human judgments. In contrast, advanced instruction-tuned models (including GPT-4o and GPT-4o-mini) achieve substantially higher positive correlations, suggesting they better reflect real-world moral attitudes. While scaling up model size and using instruction tuning can improve alignment with cross-cultural moral norms, challenges remain for certain topics and regions. We discuss these findings in relation to bias analysis, training data diversity, and strategies for improving the cultural sensitivity of LLMs.
Exploring Cultural Variations in Moral Judgments
with Large Language Models
Hadi Mohammadi, Efthymia Papadopoulou, Yasmeen F.S.S. Meijer, and Ayoub Bagheri Department of Methodology and Statistics, Utrecht University, Utrecht, The Netherlands h.mohammadi@uu.nl, evi.papado98@gmail.com, mijntje.meijer@live.nl ,and a.bagheri@uu.nl
1 Introduction
Over the past few years, LLMs have gained prominence in both academic and public discussions (Bender et al., 2021). Advances in model performance have made LLMs appealing for diverse applications, such as social media content moderation, chatbots, content creation, real-time translation, search engines, recommendation systems, and automated decision-making. While modern LLMs (e.g., GPT-4) show strong performance, a critical concern is how these models may inherit biases, including gender, racial, or cultural biases, from their training data. LLMs can easily absorb such biases because they learn from large-scale text corpora containing entrenched stereotypes (Stańczak and Augenstein, 2021; Karpouzis, 2024).
These biases raise concerns about fairness, particularly in contexts requiring moral judgments. If an LLM is trained mostly on data that negatively or inaccurately portrays certain cultural groups, it may repeat that bias in its responses. As these models become more widespread and globally deployed, the risk of perpetuating cultural biases grows, especially when moral perspectives are different from common rules or what surveys usually show. In fact, recent research shows that current LLMs often exhibit a default Western-centric bias (Adilazuarda et al., 2024), underscoring the need to evaluate their cross-cultural validity
It is crucial to see whether LLMs accurately mirror the moral judgments observed across diverse cultures. Despite its importance, this issue has received limited attention (Arora et al., 2023; Liu et al., 2024). Our study investigates whether both monolingual and multilingual Pre-trained Language Models (PLMs) can capture nuanced cultural norms. These norms include subtle ethical differences across regions, for example, the acceptance of alcohol consumption or differing attitudes on topics like abortion. Although recent research suggests that multilingual PLMs might capture broader cultural nuances, they often fall short of reflecting the moral subtleties present in less dominant cultural groups (Hämmerl et al., 2022; Papadopoulou et al., 2024).
We examine this question using two well-known cross-cultural datasets: the World Values Survey (WVS) (Inglehart et al., 2014; Haerpfer et al., 2022), and the PEW Research Center’s Global Attitudes Survey, which includes a module on moral issues across many countries (Pew Research Center, 2023). These surveys offer a detailed view of moral and cultural norms globally, serving as a benchmark for comparing LLMs outputs against actual human responses. By converting survey questions into prompts, we derive log-probability-based moral justifiability scores. We then compare these scores with survey-based consensus on various ethical issues (e.g., drinking alcohol, sex before marriage, abortion, homosexuality), allowing us to see how closely different model types and training approaches align with cultural norms. Evaluating how effectively LLMs represent cultural values has both scholarly and practical significance. If a model systematically misrepresents or overlooks certain moral perspectives, it may reinforce stereotypes or lead to biased outcomes. On the other hand, more culturally aware models can highlight both shared values and nuanced disagreements, potentially contributing to more balanced dialogue. By comparing model outputs to reliable survey data, we identify areas where LLMs align with human values and highlight gaps in capturing diverse moral perspectives.
Our contributions are threefold: (1) We introduce a structured probing framework that leverages carefully designed prompts, contrasting moral statements, and log-probability-based scoring to assess how LLMs assign justifiability values to morally complex scenarios across cultures. (2) We empirically analyze the alignment between LLM-derived moral scores and human survey responses using correlation and clustering, highlighting where models reflect or deviate from real-world moral judgments. (3) We extend our evaluation to state-of-the-art instruction-tuned and large-scale models, examining whether instruction tuning and scaling enhance alignment with cross-cultural moral norms. By identifying key strengths, weaknesses, and factors influencing model-human agreement, our work contributes to improving training data strategies, mitigating biases, and fostering the development of culturally aware language models.
2 Literature review
LLMs inherit biases embedded in their training data, and these biases can be amplified upon large-scale deployment. Because the underlying corpora often reflect entrenched social hierarchies, models run the risk of reproducing or even intensifying unfair patterns. Recent work has underscored this from multiple perspectives, a 2025 study introduced a unified framework for transparency, fairness, and privacy in AI pipelines (Radanliev, 2025), while an interdisciplinary survey emphasized the importance of diversity, equity, and inclusion (DEI) as prerequisites for trustworthy AI (Cachat-Rosset and Klarsfeld, 2023a). Taken together with earlier warnings about opaque language-model behaviors (Bender et al., 2021), these findings illustrate the need for technical innovation to proceed hand-in-hand with social safeguards. In addition to high-level ethical governance, researchers are exploring concrete mitigation strategies. For example, LLM data augmentation has improved intent-classification accuracy without sacrificing fairness, provided that the augmentation is carefully curated (Benayas et al., 2024). Complementary work on adapter tuning for non-English LLMs shows that modest architectural modifications can substantially boost performance in culturally diverse benchmarks, thereby supporting more inclusive NLP systems (Zhou et al., 2024).
Moral judgments themselves—evaluations of actions, intentions, or individuals as acceptable or objectionable—can differ widely by culture, shaped by religious traditions, social norms, and historical contexts (Haidt, 2001; Shweder et al., 1997). Understanding how such pluralistic values are (or are not) embedded in contemporary LLMs remains a pressing research concern. As noted by Graham et al. (2016), Western, Educated, Industrialized, Rich, and Democratic (W.E.I.R.D.) societies emphasize individual rights and autonomy, while non-W.E.I.R.D. societies often stress communal responsibilities and spiritual considerations. Consequently, people in W.E.I.R.D. cultures may view personal choices like sexual behavior as an individual right, while those in non-W.E.I.R.D. cultures consider them a collective moral concern. Although many moral values overlap across cultures, there are also areas of genuine divergence, often referred to as moral value pluralism (Johnson et al., 2022; Benkler et al., 2023). However, Kharchenko et al. (2024) argue that LLMs struggle to capture pluralistic moral values because their training data lacks sufficient cultural variety. Likewise, Du et al. (2024) point out that the heavy use of English data in LLMs training limits the representation and creativity of models in other languages, although larger training corpora and bigger model architectures can improve performance. Arora et al. (2023) suggest that multilingual LLMs could learn cultural values by incorporating multilingual data in their training. Yet, the limited diversity within multilingual corpora can still cause these models to perform inconsistently across languages and cultural contexts. Benkler et al. (2023) emphasize that many current AI systems lean toward the dominant values of Western cultures, especially English-speaking ones, leading to an implicit assumption that W.E.I.R.D. values are universal.
During training, LLMs use word embeddings to learn semantic and syntactic relationships based on how frequently words co-occur. These embeddings can encode the same social biases found in the training data (Nemani et al., 2024; Mohammadi et al., 2025). This association-based learning can produce biased outputs that influence the model’s fairness and reliability. For instance,  Johnson et al. (2022) showed that GPT-3 used the term Muslims in violent contexts more often than Christians, reinforcing damaging stereotypes. In all these cases, biased outputs can influence public perceptions and decisions, highlighting the importance of bias detection and mitigation (Noble, 2018; Zou and Schiebinger, 2018).
Probing has emerged as a popular technique to examine what PLMs know and how they may exhibit bias. Ousidhoum et al. (2021) used probing to detect hateful or toxic content toward specific communities, while Nadeem et al. (2021) used context-based association tests to investigate stereotypes. Arora et al. (2023) adapted cross-cultural survey questions into prompts to test multilingual PLMs in 13 languages, discovering that these models often failed to match the moral values embedded in their training languages. Although there are multiple probing approaches, from cloze-style tasks to pseudo-log-likelihood scoring (Nadeem et al., 2021; Salazar et al., 2019), each has limitations. A simpler method directly computes the probability of specific tokens, following the original transformer design (Vaswani et al., 2017).
Research on AI ethics underscores the need for models that respect cultural distinctions and support equitable treatment (Zowghi and da Rimini, 2023; Cachat-Rosset and Klarsfeld, 2023b; Karpouzis, 2024; Meijer et al., 2024). Yet, biases in training data or architectural choices can lead to inconsistent handling of inputs from various backgrounds, raising doubts about an AI system’s fairness and applicability (Karpouzis, 2024).While studies like Arora et al. (2023) and Benkler et al. (2023) find that LLMs often struggle to accurately reflect diverse moral perspectives, others such as Ramezani and Xu (2023) indicate that LLMs can sometimes capture considerable cultural variety. Similarly, Cao et al. (2023) showed that ChatGPT aligns strongly with American cultural norms while adapting less effectively to others, reinforcing concerns of Western-centric bias in LLM outputs. This discrepancy highlights the need for more research on how LLMs learn and represent moral values in different cultural settings. Even though LLMs can inherit some cultural biases, the extent of their cross-cultural fidelity remains an open question (Caliskan et al., 2017).
3 Materials and methods
3.1 Data
To evaluate cross-cultural moral attitudes, we use two datasets: World Values Survey (WVS) Wave 7 and the PEW Research Center Global Attitudes Survey 2013. Each dataset’s moral questions are labeled with topic codes. See Table 4 in Appendix A for a full reference.
World Values Survey Wave 7
The WVS conducted from 2017 to 2020111https://www.worldvaluessurvey.org/WVSDocumentationWV7.jsp, which covers respondents from 55 countries (Inglehart et al., 2014; Haerpfer et al., 2022). We use the section of the survey dealing with Ethical Values and Norms. In this section, participants were asked to rate the justifiability of 19 different behaviors or issues with moral connotations. These include topics such as divorce, euthanasia, political violence, cheating on taxes, and others. We performed preprocessing by filtering the dataset to retain only the responses to the 19 moral questions (Q177 to Q195) and the country code for each respondent.
Each response is an integer from 1 to 10. We then mapped the country codes to country names (using the provided codebook) so that each respondent entry includes their country and their answers to the moral questions. Next, we handled missing or non-response values. Entries coded as $-1$, $-2$, $-4$, or $-5$ (i.e., Don’t know, No answer, Not asked, and Missing) were set to 00, so they would not distort later calculations. We then grouped the data by country and averaged the responses for each moral statement. This yields a country-level average moral approval score for each of the 19 issues. Because different countries may use the $1$-–$10$ scale differently (culturally, some may avoid extreme ratings, etc.), and to facilitate comparison with the second dataset, we normalized these country mean scores to a range of $[-1,1]$, with $-1$ denoting never justifiable and $+1$ denoting always justifiable.
After these steps, the WVS data provides, for each country and each moral topic, a score between -1 and 1 representing how acceptable that behavior is on average according to that country’s respondents. Higher scores mean the society tends to view the behavior as more acceptable or justifiable, whereas lower scores mean it is seen as less acceptable or not justifiable. We treat these normalized country-by-topic scores as the empirical ground truth of moral attitudes.
Figure 1: Spread of responses (country mean scores) across the moral topics for the WVS Wave 7 dataset.
Figure 1 shows the spread of responses across different moral topics and countries. In other words, for each moral topic, how varied are the country scores? Some topics might have very similar scores in every culture (indicating global agreement), while others show a wide range (indicating high cross-cultural controversy).
PEW Global Attitudes Survey 2013
The PEW collected responses on moral issues from 39 countries, with about 100 respondents per country for the relevant questions [^2] https://www.pewresearch.org/dataset/spring-2013-survey-data/. Unlike WVS, which used a 10-point scale, the PEW survey questions were simpler: for each issue, respondents were asked whether the behavior is morally acceptable, morally unacceptable, or not a moral issue.
From the PEW dataset, we extracted the questions corresponding to those eight moral topics (Q84A to Q84H). We again retained only the country identifier and these responses for our purposes. We coded the responses in a numeric way to be analogous to the WVS scale: for each question, we assigned a value of $+1$ to morally acceptable, $-1$ to morally unacceptable, and 00 to not a moral issue and all non-responses (including Depends on situation, Refused, and Don’t know). As with WVS, we grouped responses by country, averaged them for each topic, and normalized the averages to $[-1,1]$. Figure 2 shows the normalized PEW values across the eight moral questions. The comparison of normalized scores for WVS and PEW by country is also presented in Appendix B, Figure 8.
Figure 2: Spread of responses across the moral topics and countries for the PEW 2013 dataset.
3.2 Methodology
Our evaluation of LLMs involves generating moral judgment scores from the models and comparing them with the two survey data. We first outline the LLMs we selected for testing, then describe how we prompted the models to obtain moral scores for each country and topic. Finally, we detail the three evaluation methods (correlation analysis, cluster alignment analysis, and models’ error analysis) that we applied to quantify the models’ performance.
Model Selection
We evaluated a broad range of transformer-based, decoder-only language models for their capacity to reflect cross-cultural moral judgments in the WVS and PEW data. Our initial set included the GPT-2 family (GPT2-B, GPT2-M, GPT2-L(Radford et al., 2019) for its coherent text generation at modest scales, as well as OPT-125 and OPT-350 (Zhang et al., 2022) to examine mid-sized behavior on ethically sensitive content. For multilingual coverage, we tested BloomZ (Muennighoff et al., 2023), Qwen-0.5, and Qwen-72 (Bai et al., 2023), aiming to see whether broader linguistic training influences moral alignment. We then studied whether larger parameter sizes or instruction tuning could improve consistency by including Gemma-9 (Mesnard et al., 2024), Llama3-8B, Llama3.3-70I (Touvron et al., 2023a), and Llama2-70 (Touvron et al., 2023b). Additional instruction-tuned models, such as Dbrx-inst (Conover et al., 2023b), MPT-30 (MosaicML, 2023), Falcon3-7B, Falcon-40I (Almazrouei et al., 2023), GPT-NeoX20 (Black et al., 2022), T5-L (Raffel et al., 2020), and Dolly-12 (Conover et al., 2023a), covered diverse training setups and parameter scales. We further compared Bloom (Scao et al., 2022) and BloomZ (Muennighoff et al., 2023) to see how instruction-specific methods affect moral responses. Finally, we examined chat-focused or proprietary systems like GPT3.5 (Brown et al., 2020), GPT4o (OpenAI, 2024b), and GPT4o-mini (OpenAI, 2024a) to determine how interactive or closed-source models handle moral content. Importantly, none of these models received additional fine-tuning on moral or cultural data, meaning our findings show their unmodified views on ethical prompts.
Prompt Structuring
Following the template-based probing framework introduced by Ramezani and Xu (2023) for fine-grained moral-norm inference across cultures, we adapt and extend their basic method to probe a wider range of LLMs. To query the models about moral judgments, we designed prompts that mimic the structure of statements about morality in different countries. Our goal was to get the model to complete statements in a way that reveals whether it thinks a behaviour is viewed as moral or immoral in a given culture. We used two main prompt templates for each country–topic pair:
P1: In $\{country\}$, $\{topic\}$ is $\{judgment\}$.
P2: People in $\{country\}$ believe $\{topic\}$ is $\{judgment\}$.
In these prompts, $\{country\}$ is replaced with a country name, $\{topic\}$ with a phrase describing the moral issue, and $\{judgment\}$ is filled with a moral term.
Moral Judgment Scores
We compute a moral score from the model for each country-topic. Let $\mathcal{L}$ be a language model. For each moral topic (e.g., drinking alcohol), we create two versions of a prompt: $M^{\text{moral}}$ and $M^{\text{nonmoral}}$. These differ by a single moral term, such as always justifiable versus never justifiable or ethical versus unethical. We then obtain $logp\bigl{(}M^{\text{moral}}\bigr{)}\quad\text{and}\quad logp\bigl{(}M^{\text{nonmoral}}\bigr{)},$ which represent $\mathcal{L}$’s tendency toward each stance. To reduce the impact of specific word choices, we repeat this process with five moral-adjective pairs  [^3] Always justifiable vs. never justifiable, right vs. wrong, morally good vs. morally bad, ethically right vs. ethically wrong, and ethical vs. unethical and compute the average difference in log probabilities: $\Delta\;=\;\log p\bigl{(}M^{\text{moral}}\bigr{)}\;-\;\log p\bigl{(}M^{\text{nonmoral}}\bigr{)}.$
We apply min–max normalization to $\Delta$ across all topics and countries, mapping $\Delta$ into $[-1,+1]$:
$\Delta_{\text{norm}}\;=\;2\,\frac{\Delta-\Delta_{\min}}{\Delta_{\max}-\Delta_{\min}}\;-\;1.$
The result is a model-based moral justifiability score $s_{i}\in[-1,+1]$. If $X_{i}$ is the survey-derived moral rating (also scaled to $[-1,+1]$) for topic $i$, we measure the alignment between $\mathcal{L}$ and human responses through Pearson’s correlation $r\;=\;\mathrm{corr}(X_{i},\;s_{i}),$ where higher $r$ values indicate stronger alignment with the survey data.
Direct Numerical Rating
In contrast, for proprietary chat models (e.g., GPT4o and GPT4o-mini), the OpenAI ChatCompletion API does not provide access to token-level log probabilities. Instead, we adopt a direct elicitation approach. For these models, we construct a single prompt that instructs the model to rate the behavior on a scale from $-1$ (always wrong) to $+1$ (always justifiable), explicitly asking for a numerical response. Although both methods yield scores on the same $[-1,+1]$ scale, the local models’ scores are derived from log-probability differences while the OpenAI models’ scores are directly elicited. Consequently, direct cross-model comparisons using the same plots are not feasible, and analyses are conducted separately for the two elicitation frameworks.
Cross-Country Correlations and Clustering
We compare each model’s cross-country correlations on a given topic to the survey-based scores. This correlation analysis shows whether a model senses that certain issues polarize particular cultures. In addition, we represent each country as a vector of moral justifiability scores and apply clustering metrics (e.g., Adjusted Rand Index or Adjusted Mutual Information) to see if a model’s country clusters match survey-derived groupings.
Comparative Prompts
We explicitly ask the model to compare two countries’ moral judgments on a given topic. We use a direct comparative prompt of the form:
Regarding the morality of $\{topic\}$, $\{countryX\}$ and $\{countryY\}$ are similar.
This tests whether the model recognizes that some pairs of countries hold similar moral views on certain topics. Overall, our pipeline of constructing moral descriptors, calculating log-probability differences, and normalizing them gives a quantitative measure of how well each language model agrees with cross-cultural moral data.
4 Results
4.1 Correlation Analysis
Pearson correlations
We first evaluated how well each model’s predicted log-prob differences align with the WVS and PEW survey scores by computing Pearson correlations ($r$). Table 1 shows the correlations for all models alongside parameter counts and significance levels.
Table 1: Pearson correlations ($r$) for WVS and PEW. Asterisks denote significance levels: * $(p<0.05)$, ** $(p<0.01)$, *** $(p<0.001)$.
Model Params WVS PEW
$r$ $p$-value $r$ $p$-value
GPT2-B 117M 0.210 *** 0.163 **
GPT2-M 355M 0.161 *** -0.094
GPT2-L 774M 0.007 -0.256 ***
OPT-125 125M 0.016 0.127 *
OPT-350 350M -0.156 *** -0.334 ***
BloomZ 560M NaN 0.443 ***
Qwen-0.5 500M -0.408 *** 0.029
Qwen-72 72B -0.078 * -0.060
Gemma-9 9B 0.440 *** 0.573 ***
Llama3-8B 8B 0.161 *** 0.151 **
Llama3.3-70I 70B 0.036 -0.038
Llama2-70 70B -0.329 *** -0.602 ***
Falcon3-7B 7B -0.312 *** -0.415 ***
Falcon-40I 40B 0.385 *** 0.671 ***
GPT-NeoX20 20B -0.078 * 0.001
Dolly-12 12B -0.247 *** 0.010
Bloom 176B -0.048 N/A
GPT3.5 0.543 *** 0.566 ***
GPT4o 0.504 *** 0.618 ***
GPT4o-mini 0.472 *** 0.678 ***
Models such as GPT4o and GPT4o-mini achieve positive correlations on both WVS and PEW, while others (e.g., Qwen-0.5, Llama2-70) yield negative correlations. Medium-scale instruction-tuned models (e.g., Gemma-9) also show moderate-to-strong alignment, indicating that training approaches and parameter size both influence agreement with survey data.
Country-Level Correlations
Next, we computed per-country correlations to see how models fare in different regional contexts. Let $\mathbf{m}_{i}$ be the vector of a model’s predicted moral scores for country $i$ across all topics, and let $\mathbf{s}_{i}$ be the corresponding vector of survey-based scores. We compute $r_{i}=\mathrm{corr}\bigl{(}\mathbf{m}_{i},\;\mathbf{s}_{i}\bigr{)}$ for each country $i$. Figure 3 shows heatmaps for WVS and PEW datasets, where each row is a model and each column is a country.
(a) WVS
(b) PEW
Figure 3: Per-country correlations, with each cell showing $r$ for a model/country. Red implies higher positive correlation, blue implies negative correlation.
In Figure 3(a), models like Gemma-9 have strong positive correlations (red squares) with local moral views across many countries, In contrast, some large-scale Llama variants exhibit negative or near-zero correlations (blue or pale squares), indicating disagreements with respondents on specific moral issues. In Figure 3(b), no model consistently performs well across all countries. For instance, Falcon-40I has strong support in parts of the Middle East, while others show areas of divergence with surveyed populations. This highlights each model’s unique strengths and weaknesses in understanding cross-cultural diversity.
Pairwise Models’ Correlations
We then examined the relationships between models by correlating their log-probability difference vectors across all country–topic pairs. For any two models $X$ and $Y$, let $\mathbf{x}$ and $\mathbf{y}$ denote their respective log-prob difference scores. We compute $\rho_{X,Y}\;=\;\mathrm{corr}\bigl{(}\mathbf{x},\;\mathbf{y}\bigr{)}$, thereby producing a pairwise correlation matrix among all models. Figure 4 shows pairwise correlations for WVS and PEW datasets. Red indicates strong similarity, while blue indicates divergence.
Figure 4(a) shows that GPT2 variants (GPT2-B, GPT2-M, GPT2-L) cluster together, indicating consistent log-probability differences within the same family. In contrast, Qwen-0.5 and Qwen-72 exhibit weak or negative correlations with instruction-tuned models like Falcon-40I and Gemma-9, suggesting a different approach to morally charged prompts. Similarly, BloomZ aligns more closely with some Llama variants than with Dolly-12 or GPT-NeoX20, reflecting differences in training methods. Figure 4(b) further reveals moderate to high correlations among related models, with GPT3.5 and GPT4o showing strong alignment, while models like Llama2-70 and Llama3.3-70I may diverge from older ones like GPT2-B. These findings highlight that instruction tuning and scale produce distinct moral stance patterns, guiding model selection for tasks requiring consistent or diverse moral reasoning and helping identify outlier models with unique stances.
(a) WVS
(b) PEW
Figure 4: Pairwise correlation heatmaps of log-prob differences for (a) WVS and (b) PEW.
4.2 Cluster Alignment
[…]
In Figure 5(a), models like GPT2-Large and GPT2 are closely grouped, with GPT2-Medium merging slightly higher. A second cluster includes Bloom, OPT-125, and Llama3-8B, showing some shared correlation. Meanwhile, Qwen-0.5, Qwen-72, and dolly-v2-12b form another moderate distance group, while large-scale or instruction-tuned models (e.g., GPT3.5-turbo, GPT4o, Falcon-40I) merge only at the top, suggesting limited similarity in their log-probability difference vectors. Figure 5(b) shows a similar structure, with some clusters differing based on the models’ responses to the morally focused PEW prompts. Notably, GPT2 and Gemma-9 cluster at low linkage heights, indicating strong similarities in their probability assignments for morally charged statements. Another cluster includes Llama2-70, Falcon3-7B, and GPT-NeoX20, which may reflect shared training data or architectural features leading to comparable moral stances.
(a) WVS
(b) PEW
Figure 5: Hierarchical clustering dendrogram
4.3 Models’ Error
Absolute Error
To assess each model’s deviation from human survey responses, we calculated the absolute difference for each country-topic pair as follows:
[…]
Figure 6 shows these distributions for WVS ( 6(a)) and PEW ( 6(b)), aggregated over all models.
Figure 6: Absolute error distributions across all models for (a) WVS and (b) PEW. Many errors cluster between 0.2 and 0.6, but some exceed 1.0.
In the case of WVS (see Figure 6(a)), many predictions fall within an error range of about $0.2$ to $0.6$, indicating that model outputs are often close to the average moral ratings provided by respondents. However, there is a significant tail extending beyond $1.0$, suggesting that for controversial or culturally sensitive topics, model predictions can diverge greatly from real human attitudes. A similar pattern is seen with PEW (see Figure 6(b)), where maximum errors rarely exceed $3.0$. While most country-topic pairs cluster around errors of $0.2$ to $1.0$, a notable number exceed $1.5$ or $2.0$, highlighting systematic misalignments in specific ethical domains that may vary widely across cultures or lack adequate representation in the training data.
Mean Absolute Error
While correlation captures how well each model’s normalized outputs align with survey responses, we also examine the Mean Absolute Error (MAE) per (model, topic) pair. This highlights which moral topics each model finds “harder” (higher error) or “easier” (lower error). Figure 7 displays a heatmap across models (columns) and topics (rows) with darker cells indicating higher error, and Tables 2 and 3 show the ten easiest and hardest topics, respectively, based on average error.
In Figure 7, topics like political violence, suicide, and stealing property result in high errors for multiple models, while issues such as drinking alcohol, using contraceptives, and divorce are generally easier for systems to manage.
In Table 2, the topic using contraceptives has the highest average error, recorded at 0.51, while the topic death penalty has a lower average error of 0.36. A low standard deviation indicates consistent ease across different models, whereas a high standard deviation suggests that only some models find the topic easy to address. In contrast, Table 3 highlights that political violence leads the list with an average error of 0.95. This is followed by suicide, stealing property, and accepting a bribe while on duty.
Figure 7: Heatmap of mean absolute errors by topic (rows) and model (columns).
Table 2: Ten easiest topics (lowest mean absolute error).
Topic Avg. Error Std. Dev.
using contraceptives 0.5111 0.2109
gambling 0.4911 0.1632
drinking alcohol 0.4815 0.1115
parents beating children 0.4622 0.2617
getting a divorce 0.4311 0.0824
having casual sex 0.4075 0.2079
divorce 0.3913 0.0723
claiming govt. benefits not entitled 0.3862 0.1991
euthanasia 0.3838 0.0792
death penalty 0.3633 0.1472
Table 3: Ten hardest topics (highest mean absolute error).
Topic Avg. Error Std. Dev.
political violence 0.9546 0.3650
suicide 0.9229 0.2486
stealing property 0.8393 0.3416
someone accepting a bribe 0.7998 0.3738
for a man to beat his wife 0.7819 0.2878
cheating on taxes 0.7170 0.3617
violence against other people 0.7091 0.3323
terrorism (political/ideological) 0.6919 0.2806
homosexuality 0.6056 0.1665
abortion 0.5985 0.3104
5 Discussion and Conclusion
Our findings show that language models vary considerably in how well they replicate cross-cultural moral judgments, as captured in the WVS and PEW surveys. Larger or instruction-tuned models, such as Falcon-40I, Gemma-9, and GPT4o, frequently demonstrate higher correlations with aggregated human survey responses. In contrast, some models, including Qwen-0.5 and Llama2-70, yield systematically negative correlations, suggesting that scale alone does not guarantee alignment with moral attitudes if the underlying training data or methodology is insufficiently diverse or biased.
In addition, topic-level analysis reveals that certain issues (e.g., political violence, terrorism, or wife-beating) consistently produce higher mean errors across different architectures. These discrepancies suggest that moral questions involving violence or extreme social norms may pose particular challenges for current language models, especially when training data do not include nuanced representations of such topics. Even models that perform relatively well on broad measures sometimes fail on region-specific or contentious issues. This trend aligns with evidence that LLMs handle clear-cut moral scenarios well but often display uncertainty or divergence on morally ambiguous dilemmas (Scherrer et al., 2023). Per-country heatmaps similarly highlight that no single model excels in all areas: while a model may align with opinions in Western nations, it can deviate markedly in communities whose moral or cultural practices are underrepresented in its training corpora.
Despite these limitations, instruction-tuned and larger models show promise in better reflecting overall moral consensus in many cases. This suggests that scaling models and using tailored training, where instructions or datasets capture diverse viewpoints, can improve moral judgment alignment. However, performance still varies, highlighting the need to analyze results in detail (e.g., by topic or country) rather than relying on a single global metric. From an applied perspective, these insights can guide the development of more culturally responsive AI systems, for example, informing content moderation policies or chatbot designs that respect regional norms.”
5.1 Limitations
Although our methodology offers insights into cross-cultural moral alignment in language models, it has several limitations that should be acknowledged. First, the WVS and PEW data capture broad national averages and may not fully reflect within-country heterogeneity, especially in regions with significant cultural or linguistic diversity. Second, our log-probability difference calculation relies on short prompt templates, which might not elicit the full context required for more complex moral issues. Third, the models we evaluated differ in size, instruction tuning, and training data composition, making it challenging to isolate the effect of each factor.
A further limitation arises from the necessity of employing distinct evaluation strategies. For local models, we have access to token-level log probabilities, enabling us to compute log-probability differences as a proxy for moral judgment. However, for OpenAI’s proprietary chat models, we rely on directly elicited numerical scores because the API does not expose internal log probabilities. This divergence means that the resulting moral scores are derived from different underlying mechanisms, precluding a direct, unified comparison of model outputs in our visualizations. Future work might seek alternative methods to bridge this gap or develop metrics that are comparable across elicitation approaches.
6 Conclusion
In conclusion, our analysis of moral stance alignment across WVS and PEW data underscores both the progress and the continuing gaps in LLMs’ performance. Models with substantial parameter counts and instruction-tuned frameworks frequently achieve moderate-to-high correlations with surveyed human judgments, suggesting an ability to capture broad moral viewpoints. However, sizable deviations persist on sensitive topics and in particular cultural contexts, indicating that no current model entirely overcomes biases or data deficiencies. Thus, while larger or more specialized training procedures can improve a model’s capacity to reflect human moral attitudes, they do not guarantee universal alignment. Future work must address these persistent shortcomings through expanded training corpora, targeted bias mitigation, and refined evaluation protocols that account for cultural and topic-level nuances.
Ethical considerations
Using language models in real-world applications has important ethical implications and risks. Even though these models can approximate broad moral opinions, they may misrepresent local or minority viewpoints if their training data is not diverse enough. This misrepresentation can lead to biases or stereotypes, especially on sensitive topics like domestic violence, religious norms, or political extremism. If a model’s output is mistakenly viewed as a true reflection of public opinion, automated decisions could unfairly target or exclude certain groups, worsening existing inequalities. Moreover, significant misalignment on controversial topics can undermine public trust if model predictions seem harmful or insensitive. To reduce such risks, it is vital to include diverse voices and expert feedback when building and testing these models. Adding regular evaluations on moral or cultural issues, transparent reports of known biases, and human review for high-stakes decisions, can help ensure ethical and responsible deployment. As language models evolve, balancing technical progress with careful oversight will be essential for maintaining fairness and trust in automated systems.
Funding
This research received no external funding.
Disclosure statement
The authors declare no conflict of interest.
Data and code availability
The full source code, experiment scripts, and processed datasets are openly available on GitHub. [^4] https://github.com/mohammadi-hadi/cultural-moral-judgments-llms
Author Contributions
H.M. and A.B. conceptualized the research. H.M., E.P., Y.M., and A.B. developed the methodology. H.M., E.P., and Y.M. contributed to software implementation, while H.M. and A.B. handled validation. E.P., Y.M., and H.M. composed the original draft, and H.M. and A.B. oversaw review and editing.
Acknowledgments
We appreciate the maintainers of WVS and PEW data for enabling large-scale cross-cultural analysis. We also thank Dr. Anastasia Giachanou for her valuable comments and feedback and Tina Shahedi for her editorial contributions to this paper. We also thank SURF for providing the computational facilities.
Appendix A Topic Codes for WVS and PEW
Table 4: Mapping of Topic Codes to the Dataset (WVS or PEW) and their corresponding moral questions.
Topic Code Dataset Moral Question
Q177 WVS Claiming government benefits to which you are not entitled
Q178 WVS Avoiding a fare on public transport
Q179 WVS Stealing property
Q180 WVS Cheating on taxes
Q181 WVS Someone accepting a bribe in the course of their duties
Q182 WVS Homosexuality
Q183 WVS Prostitution
Q184 WVS Abortion
Q185 WVS Divorce
Q186 WVS Sex before marriage
Q187 WVS Suicide
Q188 WVS Euthanasia
Q189 WVS For a man to beat his wife
Q190 WVS Parents beating children
Q191 WVS Violence against other people
Q192 WVS Terrorism as a political, ideological or religious mean
Q193 WVS Having casual sex
Q194 WVS Political violence
Q195 WVS Death penalty
Q84A PEW Using contraceptives
Q84B PEW Getting a divorce
Q84C PEW Having an abortion
Q84D PEW Homosexuality
Q84E PEW Drinking alcohol
Q84F PEW Married people having an affair
Q84G PEW Gambling
Q84H PEW Sex between unmarried adults
Appendix B WVS & PEW scores by country
Figure 8 compares normalized WVS (orange) and PEW (gold) scores by country. Each box shows the interquartile range, with medians as horizontal lines and diamonds marking outliers. The broader spread in the WVS data for many countries suggests higher variance in moral acceptance. Some countries, such as the United States or Czech Republic, show very wide ranges, from near $-1$ (never justifiable) to close to $+1$ (always justifiable). Others, often in the Middle East or South Asia, have more negative medians, reflecting stricter cultural norms on certain issues.
Figure 8: Distribution of normalized WVS (orange) and PEW (gold) survey scores by country.
Appendix C Individual Figures by Model & Dataset
In each scatter plot, the horizontal axis survey_score corresponds to WVS in Figure 9 and PEW ratings in Figure 10. Meanwhile, the vertical axis log_prob_diff shows the difference between the log-probability the model assigns to a morally justifiable statement vs. a morally unjustifiable statement. A positive slope suggests that higher survey acceptance correlates with higher log-prob differences in the same direction, meaning better alignment. Conversely, negative slopes may show systematic misalignment on that dimension.
GPT2-B
GPT2-M
GPT2-L
OPT-125
OPT-350
BloomZ
Qwen-0.5
Qwen-72
Llama3-8B
Llama3.3-70I
Llama3.3-70I
Falcon3-7B
Falcon-40I
GPT-NeoX20
Dolly-12
Bloom
Llama2-70
Figure 9: Scatter plots for WVS dataset
GPT2-B
GPT2-M
GPT2-L
OPT-125
OPT-350
BloomZ
Qwen-0.5
Qwen-72
Llama3-8B
Llama3.3-70I
Llama3.3-70I
Falcon3-7B
Falcon-40I
GPT-NeoX20
Dolly-12
Bloom
Llama2-70
Figure 10: Scatter plots for PEW dataset